OpenStudy (anonymous):

Solve and check the equation. (4x + 1)1/3 = -3 A) 2 B) -7 C) -31 D) -27/4

OpenStudy (anonymous):

distribute for the top (4x+1)1/3 \[\frac{4x+1}{3}=-3\] then you multiply three on both sides. Three's cancel on the left. which leaves you with 4x+1=-9 -1 -1 __________________ 4x=-10 \[\frac{4x}{4}=\frac{-10}{4}\]\] -5/2 none of those answers are correct

OpenStudy (anonymous):

oh yeah one of is correct i think you skipped a step

OpenStudy (anonymous):

no you have to turn 1/3 upside down and distribute you did it wrong

OpenStudy (anonymous):

Oh sorry!!